OpenStudy (anonymous):

A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air?

OpenStudy (anonymous):

I just dont understand how to solve.

OpenStudy (anonymous):

let x = speed of plane in still air (x+40) = speed of plane downwind (x-40) =speed of plane against the wind distance = speed *travel time

OpenStudy (anonymous):

downwind distance = headwind distance 5(x+40) = 6(x-40) 5x+200=6x-240 6x-5x=240+200 x=440 mph ans: The speed of the plane in still air is 440 mph